OpenStudy (anonymous):

Evaluate for a = 4 and b = -6. Select the simplest form. a/2b Please Explain how you got the answer so I can successfully complete the rest (: Thanks! I do online and it's Christmas break and I have no one to explain this.

myininaya (myininaya):

\[\frac{4}{2(-6)}\] a is 4 b is -6 i just plugged them in

OpenStudy (anonymous):

substitute values for a and b: a/2b = 4/(2*-6) = -4/12 Simplify by dividing both -4 and 12 by the same number.. They can both be divided by 4 so: -4/12 = (-4/4)/ (12/4) =-1/3

myininaya (myininaya):

now 4 is 2*2 does anything cancel?

myininaya (myininaya):

and then something else will cancel you can do it henkjan's way too